Startup Night 6 - 20170907 UT

Observers: JHill, BRothberg, A.Cardwelll (Tucson)
Telescope Operator: G.Bechetti (LBTO)
AO Support: D.Miller
Telescope Support: P.Hartley (LBTO)
Software Support: CBiddick, KSummers (Tucson)
Instrument Support: J.Morris (LBTO)
Engineering: McMahon

Plan

  1. New Collimation and Pointing model testing - both PEPSI polarimeters are now installed
  2. Old Version of Pointing Model (Dave T.) - requires mapping multiple stars across the sky (to compare different methods)
  3. AGW-1 Transform and AGW-2 confirmation test of Transform results
  4. Verify Field Aberrations
  5. LBC Twilight Flats
  6. LBC Spectro-photometric standard at start (BD33), photometric standard
  7. LUCI-Binocular G210 Spectroscopy test (acq, take data, take telluric, take calibrations at end of night)
  8. Other:
  • AGW3 and AGW4 PEPSI (poor seeing)
  • M3 Adjustment LBTI
  • Photometric Standards LUCI-Bino


UT 01:21 - Started up LUCIs and LBCs

UT 01:23 - Taking test biases for LBCs

LBC Sky Flats

UT 01:47 - Telescope Open - prepping for SkyFlats for LBC

UT 02:05 - LBC-B U Bessel and LBC-R V-Bessel PA=0 (first exposures 16k counts, last exposures 8k counts)

UT 02:09 - LBC-B U Bessel and LBC-R V-Bessel PA=180 (scaletime = 3) first exposures are 15k, last exposures are 7k counts

UT 02:17 - LBC-B g-sloan and LBC-R I-Bessel PA=180, scaletime =4 (first exposures 10k counts, last exposure 7k)

UT 02:21 - LBC-B g-sloan and LBC-R I-Bessel PA=0, scaletime = 7 (this might not work, first counts are 6k, last are 5k)

UT 02:29 - LBC-B g-sloan and LBC-R I-Bessel PA=0 scaletime=12 (might get some but a low of out of focus stars are present) - last bunch have lots of stars

UT 02:34 - moving to BD 33 2642 for collimation and co-pointing

UT 02:46 - dohybrid paused - couldn't find data files in /newdata, but found them eventually

UT 02:47 - doing co-pointing

UT 02:51 - moving to BD 33 2642 target field

UT 03:02 - LBC Error:

02:57:05.993000 bad input voltage line 7 (+23.90V instead of +25.0V+-1.0V) - not sure if this is related

2017/09/07 02:57:18.565167 E B >>> preset/expose/save interrupted at expose time
2017/09/07 02:57:18.666233 N B completed 6 OB image/s [src/lbc/actions.c:1159]
2017/09/07 02:57:18.666266 E B OB execution error: one or more systems failed [src/lbc/actions.c:1162]
2017/09/07 02:57:18.666394 N - freeing FITS headers
2017/09/07 02:57:18.666750 N - UI update with message >(BLUE Channel) OB execution error: (CAMERA) CCD Controller error!< at 1504753038.666523
2017/09/07 02:57:18.674867 N - UI update at 1504753038.666870
2017/09/07 02:57:18.675005 N - executing PLAY command completed
OB quit, I am reloading and restarting OB

UT 03:08 - same error OB stops at exposure 6

2017/09/07 03:06:07.481765 N R CAMERA >>> save completed
2017/09/07 03:06:07.481915 N R save completed
2017/09/07 03:06:07.481963 N R >>> preset/expose/save completed
2017/09/07 03:06:07.583054 N R completed 6 OB image/s [src/lbc/actions.c:1176]
2017/09/07 03:06:07.583088 N R OB execution step completed
2017/09/07 03:06:07.583165 N - freeing FITS headers
2017/09/07 03:06:07.583481 N - UI update with message >(BLUE Channel) OB execution error: (CAMERA) CCD Controller error!< at 1504753567.583283
2017/09/07 03:06:07.589096 N - UI update at 1504753567.583598
2017/09/07 03:06:07.589237 N - executing PLAY command completed
2017/09/07 03:06:08.432000 S R CAMERA CCDCTRL electronics temperature: 55.8 [Celsius]
2017/09/07 03:06:08.588000 N R CAMERA CCDCTRL issuing >tftp -b32768 -o 192.168.38.32 put lbcr.20170907.030525.fits d:\lbcr.20170907.030525.fits > nul 2>&1< upload command completed in 3.0 sec
2017/09/07 03:06:08.635000 S R CAMERA CCDCTRL file upload queue contains 0 file/s
2017/09/07 03:06:08.963000 S R CAMERA CCDCTRL original temperature reading value: 16922.966380
2017/09/07 03:06:08.994000 S R CAMERA CCDCTRL detectors temperature: 190.9 [Kelvin]
2017/09/07 03:06:09.026000 N R CAMERA CCDCTRL readout ...

in log:

  1. /09/07 03:03:46.510000 CAMERA CCDCTRL input voltage line 9 is +5.13V ( +4.8V+-1.0V)
    789 2017/09/07 03:03:46.510000 CAMERA CCDCTRL input voltage line 8 is -24.70V (-25.0V+-1.0V)
    788 2017/09/07 03:03:46.478000 CAMERA CCDCTRL bad input voltage line 7 (+23.90V instead of +25.0V+-1.0V)
    787 2017/09/07 03:03:46.478000 CAMERA CCDCTRL input voltage line 6 is -11.03V (-11.0V+-1.0V)
    786 2017/09/07 03:03:46.463000 CAMERA CCDCTRL input voltage line 5 is +10.75V (+11.0V+-1.0V)
    785 2017/09/07 03:03:46.463000 CAMERA CCDCTRL input voltage line 4 is -18.43V (-18.5V+-1.0V)
    784 2017/09/07 03:03:46.447000 CAMERA CCDCTRL input voltage line 3 is +18.27V (+18.5V+-1.0V)
    783 2017/09/07 03:03:46.432000 CAMERA CCDCTRL input voltage line 2 is -14.82V (-15.0V+-1.0V)
    782 2017/09/07 03:03:46.432000 CAMERA CCDCTRL input voltage line 1 is +15.18V (+15.0V+-1.0V)

Turning off Blue and testing script with Red only (should fail after 3rd exposure) - got to 5th exposure on Red no problem. stopped script. Turning off and on LBCs again see if problem persists.

UT 03:15 - turned off LBCs, will turn them back on again.

UT 03:20 - LBCs back on, sending OB again

UT 03:23 - turning off and on seemed to fix this. errant voltage bug on LBC-B?

UT 03:34 - error cropped up again at 34/48

2017/09/07 03:34:04.828938 N R CAMERA >>> save completed
2017/09/07 03:34:04.829098 N R save completed
2017/09/07 03:34:04.829147 N R >>> preset/expose/save completed
2017/09/07 03:34:04.930190 N R completed 33 OB image/s [src/lbc/actions.c:1176]
2017/09/07 03:34:04.930223 N R OB execution step completed
2017/09/07 03:34:04.930324 N - freeing FITS headers
2017/09/07 03:34:04.930626 N - UI update with message >(BLUE Channel) OB execution error: (CAMERA) RPC Client/Host communication failed!< at 1504755244.930413
2017/09/07 03:34:04.938490 N - UI update at 1504755244.930743
2017/09/07 03:34:04.938627 N - executing PLAY command completed
From LBC parsing log for Blue:

2398 2017/09/07 03:33:42.525023 OB execution error: one or more systems failed [src/lbc/actions.c:1162]
2397 2017/09/07 03:33:42.524980 completed 32 OB image/s [src/lbc/actions.c:1159]
2396 2017/09/07 03:33:42.474060 >>> preset/expose/save interrupted at preset time, retcode=7
2395 2017/09/07 03:33:42.474033 preset completed
2394 2017/09/07 03:33:42.473942 camera system encountered a problem 6 [src/channel/channel.c:1138]
2393 2017/09/07 03:33:42.473835 CAMERA RPC Client/Host communication failed (4) [src/camera/camera.c:1169]
2392 2017/09/07 03:33:42.473780 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2391 2017/09/07 03:33:40.472574 CAMERA preset failed with 4 (or NULL return) try again ...
2390 2017/09/07 03:33:40.472532 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2389 2017/09/07 03:33:38.471303 CAMERA preset failed with 4 (or NULL return) try again ...
2388 2017/09/07 03:33:38.471239 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2385 2017/09/07 03:33:36.470037 CAMERA preset failed with 4 (or NULL return) try again ...
2384 2017/09/07 03:33:36.469996 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2383 2017/09/07 03:33:34.730595 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1033]
2382 2017/09/07 03:33:34.729374 CAMERA reconnection attempt after 10 wait loops since last failure
2381 2017/09/07 03:33:34.468766 CAMERA preset failed with 4 (or NULL return) try again ...
2380 2017/09/07 03:33:34.468721 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2379 2017/09/07 03:33:32.467488 CAMERA preset failed with 4 (or NULL return) try again ...
2378 2017/09/07 03:33:32.467446 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2377 2017/09/07 03:33:30.466209 CAMERA preset failed with 4 (or NULL return) try again ...
2376 2017/09/07 03:33:30.466166 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2375 2017/09/07 03:33:28.464993 CAMERA additional error info: 104 [src/camera/camera.c:1136]
2374 2017/09/07 03:33:28.464949 CAMERA preset failed with 4 (or NULL return) try again ...
2373 2017/09/07 03:33:28.279309 CAMERA reconnection successful [src/camera/camera.c:1147]
2370 2017/09/07 03:33:26.278082 CAMERA preset failed with 4 (or NULL return) try again ...
2369 2017/09/07 03:33:26.278038 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2368 2017/09/07 03:33:24.276818 CAMERA preset failed with 4 (or NULL return) try again ...
2367 2017/09/07 03:33:24.276777 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2366 2017/09/07 03:33:22.275552 CAMERA preset failed with 4 (or NULL return) try again ...
2365 2017/09/07 03:33:22.275510 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2364 2017/09/07 03:33:20.274268 CAMERA preset failed with 4 (or NULL return) try again ...
2363 2017/09/07 03:33:20.274225 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2362 2017/09/07 03:33:18.272986 CAMERA preset failed with 4 (or NULL return) try again ...
2361 2017/09/07 03:33:18.272934 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2358 2017/09/07 03:33:16.728490 CAMERA closing connection for a while
2357 2017/09/07 03:33:16.271682 CAMERA preset failed with 4 (or NULL return) try again ...
2356 2017/09/07 03:33:16.271628 CAMERA clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused [src/camera/camera.c:1143]
2355 2017/09/07 03:33:14.778519 CAMERA RPC Client/Host communication failed [src/camera/camera.c:1019]
2354 2017/09/07 03:33:14.728487 CAMERA StatusPatch RPC error: 3 [src/camera/camera.c:1519]
2353 2017/09/07 03:33:14.270376 CAMERA additional error info: 104 [src/camera/camera.c:1136]
2352 2017/09/07 03:33:14.270326 CAMERA preset failed with 4 (or NULL return) try again ...
2351 2017/09/07 03:33:13.509000 CAMERA CCDCTRL electronics temperature: 47.1 [Celsius]

UT 03:41 - made a new OB called BD33_Srg_FzY_LastDither that does the last dither position to complete the script

UT 03:42 - GUI shows ERROR on expd images even when loading OB - has 12 exposures moves to 0.0,-635.0

This is IT 6761

UT 03:58 - LBC Blue appears to be frozen now, in power control Camera & Shutter show STOP - no warning appeared anywhere else

Powering off LBCs again, will try one more time

Powered off shows warning on LBC Blue Camera & Shutter

4024 2017/09/07 04:00:05.666513 CAMERA ccdctrlclose failed; RPC Client/Host communication failed (3) [src/camera/camera.c:665]

3936 2017/09/07 03:58:41.471846 camera system encountered a problem: RPC Client/Host communication failed (6) [src/channel/channel.c:874]

3934 2017/09/07 03:58:41.471723 CAMERA ccdctrlstop failed; RPC Client/Host communication failed (RPC error code:3) [src/camera/camera.c:842]

3781 2017/09/07 03:58:02.400569 >>> preset/expose/save interrupted at preset time, retcode=4

10 2017/09/07 03:43:16.127676 camera system encountered a problem: RPC Client/Host communication failed (6) [src/channel/channel.c:874]
9 2017/09/07 03:43:16.127544 CAMERA ccdctrlstop failed; RPC Client/Host communication failed (RPC error code:3) [src/camera/camera.c:842]
8 2017/09/07 03:43:01.544488 OB execution error: one or more systems failed [src/lbc/actions.c:1162]
7 2017/09/07 03:43:00.303606 camera system encountered a problem 6 [src/channel/channel.c:1138]
6 2017/09/07 03:43:00.303490 CAMERA RPC Client/Host communication failed (4) [src/camera/camera.c:1169]
5 2017/09/07 03:43:00.217939 camera system encountered a problem: RPC Client/Host communication failed (6) [src/channel/channel.c:874]
4 2017/09/07 03:43:00.217815 CAMERA ccdctrlstop failed; RPC Client/Host communication failed (RPC error code:3) [src/camera/camera.c:842]
3 2017/09/07 03:33:42.525023 OB execution error: one or more systems failed [src/lbc/actions.c:1162]
2 2017/09/07 03:33:42.474060 >>> preset/expose/save interrupted at preset time, retcode=7
1 2017/09/07 03:33:42.473942 camera system encountered a problem 6 [src/channel/channel.c:1138]
0 2017/09/07 03:33:42.473835 CAMERA RPC Client/Host communication failed (4) [src/camera/camera.c:1169]

UT 04:04 - Powering on LBCs again - blue came back up ok

UT 04:08 - attmepting final dither position for BD33 one last time (airmass 1.56 now)

UT 04:13 - BD33 completed (slightly less than the 48 exposures in the original OB, 44/48)

Pointing Model Data Collection

UT 04:13 - Reconfiguring for LUCI-Binocular

UT 04:42 - Running Dave's Pointing Program

- will collimate and point at first object, then run TRACK mode and manually take LUCI/LUCI exposures

UT 05:24 - sending preset to collimate and point, had issues with AGWs and tertiary. AdSec hit a limit.

UT 05:29 - PCS pointing log started

UT 05:43 - still working on pointing and collimation startup
UT 05:46 - Starting with PT10m093 Track Preset PA=0

PtStars PtStars _10m.dat 21:45:00 EL=30

26 good points found with EL>30
Sorted in Azimuth:

AZ ALT RA J2000 DEC PtStarID UT L1 File# L2 File#
======= ===================== ==== ===============
356.90 62.30 21:33:21.281 60:19:54.02 PT10m093 05:48 0194 0008

16.49 46.08 00:19:22.047 71:35:43.13 PT10m001 05:52 0195 0009 - Binary Star - should not be used
42.58 64.91 23:28:13.610 48:46:01.30 PT10m101 05:56 0196 0010
56.85 48.32 01:13:44.415 45:11:40.71 PT10m008 05:58 0197 0011
75.21 32.50 02:15:28.469 28:07:25.47 PT10m011 06:00 0198 0012
91.47 45.64 01:00:18.152 21:47:27.61 PT10m005 06:08 0199 0013 (blurry) 0014 (looks same as 0013)
94.35 69.43 23:19:22.350 28:54:26.91 PT10m100 06:10 0200 0015
117.70 55.17 23:49:52.999 12:42:56.46 PT10m102 06:18 0201 0016 -collimated on this star, pointing was way off, went back to 100, then back here for data
124.55 40.25 00:20:49.768 -0:52:15.96 PT10m002 06:20 0202 0017
156.03 31.09 23:13:20.634 -22:18:04.73 PT10m099 06:24 0203 0018
158.96 74.49 22:08:10.850 18:05:44.56 PT10m095 06:30 0204 0019
170.38 53.93 22:07:36.585 -2:58:20.23 PT10m096 06:32 0205 0020
186.32 39.76 21:24:39.566 -17:18:10.33 PT10m092 06:34 0206 0021 (ignore L2#0022)
202.98 30.22 20:18:42.543 -23:25:29.83 PT10m088 06:36 0207 ----- (GEIRS 2 error, OSA moved telescope before could recover )
206.10 62.54 20:57:45.469 7:31:09.11 PT10m091 06:39 0208 0023

202.98 30.22 20:18:42.543 -23:25:29.83 PT10m088 06:40 0209 0024
224.94 39.56 19:32:01.129 -6:36:31.88 PT10m085 06:41 0210 0025
248.70 53.32 19:24:38.008 14:31:15.84 PT10m084 06:44 0211 0026
251.39 33.43 18:15:06.740 4:12:53.32 PT10m078 06:45 0212 0027 (ignore L2 #0028)
258.08 72.25 20:26:23.657 27:29:09.41 PT10m089 06:47 -collimating here, can't see anything, sending an acquire preset to m089, AdSecs RIP'd, still having issues at UT 07:02

Finally got an active preset, collimating on star UT 07:10

258.08 72.25 20:26:23.657 27:29:09.41 PT10m089 07:13 0213 0029

281.50 50.86 18:38:33.319 31:39:38.53 PT10m081 07:14 0214 ---- (missed taking this)
309.46 61.44 19:34:12.992 46:54:10.91 PT10m086 07:17 0215 0030

281.50 34.00 18:38:33.319 31:39:38.53 PT10m081 07:17 0216 0031
355.87 83.82 21:42:43.045 38:51:56.59 PT10m094 07:20 0217 0032
356.90 62.30 21:33:21.281 60:19:54.02 PT10m093 07:21 0218 0033
199.88 74.59 22:08:10.850 18:05:44.56 PT10m095 07:24 0219 0034
358.94 73.93 23:28:13.610 48:46:01.30 PT10m101 07:26 0220 0035

380.17 83.53 BS9188 07:33 0221 0036

UT 07:35 - Finished Pointng Collection

FIELD ABERRATION DATA

Doug M. Field Data Results: -from BS9177_new
=> SX Field Data Results
             Measured bi-nodal astigmatism

                  BS9177_new
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  10.9", Delta IE = -10.9"
         Tilt = -13.4", Delta CA =  13.4"


 => DX Field Data Results
             Measured bi-nodal astigmatism

                  BS9177_new
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  24.2", Delta IE = -24.2"
         Tilt =  -7.5", Delta CA =   7.5"

THESE DELTA IE/CA need to be applied AFTER THE NEW POINTING MODEL.

Tip and Tilt from above were applied with coma-free pointing to M2 alone.

THESE MANUAL POINTING CORRECTIONS WILL BE GLOBAL OFFSETS TO THE COLLIMATION MODEL. (They were installed on 20170908 UT.)

SX: Tip = 10.9" Delta IE = -10.9" (rounding to 11) Tilt = -13.4" Delta CA = 13.4" (rounding up to 13)

SX ManualPointing X=+0.664 Y=-0.562 Z=0 RX=+108.0 RY=+127.7

DX: Tip = 24.2" Delta IE = -24.2" (rounding to 24) Tilt = -7.5" Delta CA = 7.5" (rounding to 8)

DX ManualPointing X=+0.409 Y=-1.226 Z=0 RX=+235.7 RY=+78.6

Using BS9188_short.list for Field Aberration data confirmation (8 GS in list)

UT 08:24 - SX missed GS18 - no guidestar. resent script

UT 08:25 - SX missed GS18 - run acquire preset

UT 08:28 - pointing check run binocular which killed DX - will have to redo both ignore everything up to this for aberration
UT 08:30 - need to return to BS9188

UT 08:33 - restarted Field Aberration Confirmation

UT 09:11 - DX failed to find GS7, completed 7/8 stars,

Photometric Fields for LUCI-Bino, z,J,H,K,Ks

UT 09:15 - Pointing Check for FS10

UT 09:23 - Moving to FS10, L1# 0223-0247 , L2# 0038-0062 Airmass 1.55-1.35

UT 09:27 - Keyhole and other features remain on LUCI-1 (Ks 223-227)

UT 09:32 - #0225 L1 seemed to offset out of patrol field, but for L2 it was fine

UT 09:39 - HBS alarms are going off, but can't reach OSA on phone

UT 09:41 - Error in GUI but no warning, saw something in scheduler relating to error committing setup UT 09:39:09

UT 10:12 - FS10 finished

UT 10:13 - moving NGC 205 G210 test (H-band 0.5" LS)

UT 10:14 - LUCI-2 MOS Error:

13975854 2017-09-07 10:14:14.4700 ERROR ***** Luci TWO TWO_MOSUnit Error in sequence MOS - storage to turnout. WARNING holding current still active at [card number: 1, motor number: 5, ty...
UT 10:16 - Dave T. called - no steps moved, Barry searching for more info.

  • MOS storage to turnout shows error in red in status of History of last executed sequence in LUCI-2 MOSUnit GUI - head rotation noted in message server
  • Rotator is 362 RBFG, elevation 55 degrees
  • translator was one step out of tolerance, MOS init got i back to "in turnout"
  • However, both translator and rotator holding currents are on, need to put mask back to FPU to make sure they go off (Dave T. working on that)

UT 11:05 - resending preset to target, LUCI-2 MOS ok now

UT 11:07 - SX pointing failed, DX shows star in acquisition - preset failed

UT 11:14 - new script with better guidestar - resent acquisition

UT 11:30 - offset to 2nd dither guidestars couldn't be seen, but according to OT its a valid position

UT 11:34 - Moving to Telluric star HD 9711

UT 11:42 - Taking Telluric spectra

UT 11:40 - finished Telluric

UT 11:50 - Moving back to FS10 Photometric std field - Airmass 1.2

UT 11:52 - LUCI-2 didn't switch to mirror, stayed at G210 HiRes no tilt - had to manually put in Mirror (after I paused it)
13982867 2017-09-07 11:53:55.1990 SYSTEM *** Luci TWO TWO_GratingUnit grating unit at position -1, order -, wavelength -, tilt voltage -, loop opened, tilt unstable, status: active
13982870 2017-09-07 11:53:55.1980 SYSTEM *** Luci TWO TWO_CameraUnit camera unit at position -1, status: active
LS1 #0262-0284 (Ks-z - zband not subtracting well only did 3 of them)

LS2 #0077-0099 (Ks-z - zband not subtracting well only did 3 of them)

UT 12:31 - closing up

UT 12:41 - Running LUCI 1 & 2 G210 H-band 0.5" longslit calibrations
  • Running LUCI-1 as script (285-294), LUCI-2 by hand (100-109) for Flats
  • LUCI-1 H-band flats seem a little low - about 7k vs. 9k for LUCI-2
  • LUCI-2 Grating stabilization fault, G210 shows 1micron unstable
  • Arcs, running LUCI-1 by script, LUCI-2 by hand
  • LUCI-2 (110-113, lights off, then Ar, lights off, then Xe)
  • LUCI-1 Arcs are missing dark frames, will take them manually (295 Xe, 296 Ar, 297 20sec dark, 298 60sec dark
UT 12:54 - putting LUCI-2 into safe mode

UT 12:59 - putting LUCI-1 into safe mode

AGW-1 Log:

1) UT 05:09 LFBG Rotator = 367, Elevation 75 degrees, Position Angle = -147, didn't move off-axis, was stuck on-axis

2) UT 05:34 Probe moving slowly, LFBG Rotator = 264, elevation 55 degrees, Position Angle = 0

3) UT 05:57 Probe failed to move LFBG Rotator = 127, elevation 64 degrees, Position Angle = 0

4) UT 06:36 Probe failed to move LFBG Rotator = -7.6 elevation 38 degrees, Position Angle = 0

Summary

  1. Completed Pointing Model of 24 stars (Dave T.)
  2. Completed Spectrophotometric Std BD33 2346 for LBC
  3. Completed Skyflats for LBC (LBC-B: U-Bessel 0/180, g-sloan 180; LBC-R: V-Bessel 0/180, I-Bessel 180)
  4. Completed Field Aberration Confirmation Data (Doug M.)
  5. Completed Standard Field (FS10) for LUCI-Bino z,J,H,K,Ks at two airmasses
  6. Completed G210 science-like acquisition and observation along with telluric observation
  7. Completed LUCI-Bino dithering within an imaging field (dice 5 dither)
  8. Completed G210 H-band Flats and Arcs Binocular calibrations

-- %USERSIG{BarryRothberg - 2017-09-07}%
Topic revision: r7 - 08 Sep 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback